OpenStudy (anonymous):

Find an expression for the function whose graph is the line segment joining the points (-5,10) and (7,-10).

OpenStudy (anonymous):

I was trying to use y - y1 = m (x - x1) but I could not get it to work.

OpenStudy (reemii):

it's right. m is "delta y" / "delta x". what is your "m" ?

OpenStudy (anonymous):

-5/4

OpenStudy (reemii):

yes. for the rest, choose ONE point from which you will use the coordinates as x1 and y1. don't mix using the two points.

OpenStudy (anonymous):

I calculated m wrong the first time! Thanks for the help!

OpenStudy (anonymous):

actually when I plug in 7 I do not get the correct answer..

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 10}})\quad &({\color{red}{ 7}}\quad ,&{\color{blue}{ -10}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)